OpenStudy (jamanorthi):

insert grouping symbole in the expression so that the value of the expression is 14 1. 9+39+22/11-9+3

OpenStudy (cwrw238):

(9+39+22)/(11-9+3) = 70 / 5 = 14

OpenStudy (jamanorthi):

okay

OpenStudy (jamanorthi):

thanks

OpenStudy (cwrw238):

yw